### v1

@@ 4 Practical implementation
As to the preconditioners, Chronos provides both static and adaptive pattern FSAI (sFSAI or aFSAI), which can be used as standalone preconditioners or as smoothers within a multigrid hierarchy. Chronos also provides an AMG preconditioner implementing a classical coarsening (that is, a division of the nodes into fine and coarse with no aggregation) and several interpolation schemes. Namely, classical and ExtI for Poisson-like problems and LSF for elasticity problems. Additionally, AMG’s quality is improved by using energy minimisation [32] and its application cost reduced through prolongation filtering.
@@ 5 Experimental results
The numerical experiments consist of solving Eq. 48 using three industrial test cases: the flow around a realistic car model, a finned-tube heat exchanger, and the simulation of a wind farm. While the first two allow for exploiting reflection symmetries, the last allows for studying the further advantages of translational and rotational symmetries. For simplicity, the right-hand side (RHS) used in the tests is a random vector, which ensures a broad spectrum of frequencies in the resulting residual, allowing, in turn, for a complete assessment of the preconditioner’s effectiveness in damping all error components. All the executions rely on combined MPI and multithreaded parallelism and have been conducted on the JFF cluster at the Heat and Mass Transfer Technological Center. Its non-uniform memory access (NUMA) nodes are equipped with two Intel Xeon 6230 CPUs (20 cores, 2.1 GHz, 27.5 MB L3 cache and 140 GB/s memory bandwidth) linked to 288GB of RAM and interconnected through 7 GB/s FDR Infiniband.
@@
The coarsening ratio and average nonzeros per row in Tables 5, 6, and 7 correspond to the operator at the first level of the multigrid hierarchy, i.e., after the first coarsening, which, in the case of AMGR, is induced by the reduction framework. Additionally, $n_{b}$ denotes the number of subdomains in which the domain can be decomposed, i.e., the number of repeated blocks in $\bar{K}=\mathbb{I}_{n_{b}}\otimes K$.
@@
At the top of Fig. 4, there are displayed the two test cases considered for assessing the performance of AMGR when exploiting reflection symmetries. Namely, the simulation of the flow around the DrivAer fastback car model [27] and within a finned-tube heat exchanger, on which we exploit one and three reflection symmetries, respectively. The results for the DrivAer case in Table 5 make apparent the advantages of AMGR, which preserves the excellent convergence of the standard AMG and, thanks to the faster top-level smoothing granted by replacing SpMV with SpMM, yields 43% speed-ups. The larger the number of repeated blocks in $\bar{K}=\mathbb{I}_{n_{b}}\otimes K$, the greater the benefits of replacing SpMV with SpMM. In this sense, the heat exchanger problem extends the previous results up to 3 symmetries. According to Table 6, regardless of $n_{b}$, the aggressive coarsening induced by the multigrid reduction results in almost the same coarse operator. Indeed, both the coarsening rate and density remain roughly constant. Nevertheless, larger $n_{b}$ implies faster top-level smoothing, which grants up to 68% accelerations.
@@
preconditioner $n_{b}$ coarsening ratio avg nnzr its t-sol (s) speed-up
@@
preconditioner $n_{b}$ coarsening ratio avg nnzr its t-sol (s) speed-up
@@
AMGR 2 0.14 37.5 19 1.12 1.38
AMGR 4 0.15 37.4 19 1.03 1.50
AMGR 8 0.15 37.6 18 0.91 1.68
The results in Tables 5 and 6 make clear the potential of applying AMGR on domains arising from recurring structures, such as the $6\times 4$ wind farm in Fig. 4. Indeed, while reflection symmetries are generally restricted to $n_{b}\leq 8$, rotational and translational symmetries are not, allowing, for instance, the simulation of an arbitrarily large wind farm by just discretising a single wind turbine. Table 7 summarises the results obtained on a $6\times 4$ wind farm discretised according to [15]. The wind turbines are introduced through the immersed boundary method and, therefore, discretised with a structured grid stretched around the blades and coarsened vertically towards the atmospheric boundary layer. The first thing to note is that the optimal number of subdomains into which dividing the wind farm is not $n_{b}=24$. In fact, relatively small values of $n_{b}$ led to maximum speed-ups, as the application of AMG on the reduced operator quickly counterbalanced SpMM’s accelerations. This follows from the fact that the larger $n_{b}$, the faster SpMM, but also the larger $n_{\text{ifc}}$. Then, according to Eq. 46, the less effective the multigrid reduction and, as observed in Table 7, the larger the coarsening ratio. The wind farm problem is susceptible to such a trade-off due to the reduction itself being less aggressive, as the larger coarsening ratios indicate. Nevertheless, AMGR still proved effective, being up to 70% faster than the standard AMG.
The coefficient matrices arising from the industrial CFD applications considered are remarkably sparse. As a result, so is the optimal top-level smoother, and the best results were obtained using a very light FSAI. Despite obtaining significant speed-ups, this compromised the advantages of AMGR. Not only by making SpMM’s acceleration quickly counterbalanced by the application of AMG on the reduced operator, but also by making both AMG and AMGR have comparable memory footprints. In this sense, higher-order schemes or, more generally, applications entailing a denser Poisson’s equation (e.g., linear elasticity or geomechanical problems) would strengthen the benefits of SpMM and, therefore, of AMGR.
@@
preconditioner $n_{b}$ coarsening ratio avg nnzr its t-sol (s) speed-up
@@
When it comes to the scalability of AMGR, it is worth recalling Section 4. As discussed, to replace SpMV with SpMM, it is necessary to apply a consistent domain partitioning. Then, instead of distributing the rows of $\bar{K}=\mathbb{I}_{n_{b}}\otimes K$ among all the available resources, we need to distribute the smaller $K$ so that its coefficients are effectively reused in the SpMM (analogously for the top-level smoother). Of course, distributing the smaller sub-matrix entails potentially larger communications. According to the results of Fig. 5, this is not a severe issue. On the one hand, AMGR preserves the excellent weak scalability of AMG. On the other, even if AMGR’s strong scalability is affected by the extra communication overheads, this effect is far from critical, especially considering the relatively large workloads that extreme-scale simulations typically entail.
@@ 6 Conclusions
This paper presented a multigrid reduction framework for accelerating AMG on regular domains. Given that modern supercomputers prioritise FLOP performance, most computational physics applications are memory-bound and, therefore, unable to achieve hardware’s theoretical peak performance. To address this limitation, we first showed that given an arbitrarily complex geometry presenting reflection, translational or rotational symmetries, it is possible to apply a consistent ordering that makes the coefficient matrix (and preconditioners) satisfy a regular block structure. This, in turn, allows the standard SpMV to be replaced with the more compute-intensive SpMM, accelerating the application of the preconditioners. This paper focused on accelerating AMG owing to its numerical and computational effectiveness. However, it is worth noting that the intermediate strategies presented naturally apply to other preconditioners.
@@
The proposed multigrid reduction framework introduces an aggressive coarsening to the multigrid hierarchy, reducing the memory footprint, setup and application costs of the top-level smoother. While preserving the excellent convergence of AMG, replacing SpMV with SpMM yielded significant speed-ups. Remarkably enough, AMGR, the resulting preconditioner, does not have any specific requirements on the boundary conditions, only on the mesh and its ordering. Numerical experiments on industrial CFD applications demonstrated up to a 70% speed-up in the solution of Poisson’s equation by AMGR compared to AMG. Furthermore, even if AMGR entail higher communication overheads, strong and weak scalability analyses revealed no serious degradation compared to AMG.
@@
The coefficient matrices arising from the incompressible CFD applications considered are remarkably sparse and, as a result, so is the optimal top-level smoother. Indeed, the best results were obtained smoothing with a very light FSAI and, even if we obtained significant speed-ups, this compromised the advantages of AMGR. Firstly, leveraging relatively few symmetries led to maximum speed-ups, as SpMM’s accelerations were quickly counterbalanced by the application of AMG on the reduced operator, whose size grows slightly with the number of subdomains. Secondly, both AMG and AMGR had a comparable memory footprint. Again, this follows from the particularly lightweight smoother required by the cases considered. In this sense, higher-order schemes or, more generally, applications entailing a denser Poisson’s equation (e.g., linear elasticity or geomechanical problems) would strengthen the advantages of SpMM and, therefore, AMGR.
@@
For all these, immediate lines of work include applying AMGR to (denser) problems arising from structural mechanics. Additionally, we plan to optimise the setup phase, which was irrelevant for the incompressible CFD applications considered but is crucial for simulations with variable Poisson’s equation. Significant accelerations are expected from computing the top-level smoother on the base mesh instead of the entire domain. Finally, we want to tackle the GPU implementation of AMGR and study the extension of the multigrid reduction framework towards nonregular domains, which, despite not exploiting SpMM, would benefit from the aggressive coarsening.

### v2

@@ 4 Practical implementation
As to the preconditioners, Chronos provides both static and adaptive pattern FSAI (sFSAI or aFSAI), which can be used as standalone preconditioners or as smoothers within a multigrid hierarchy. Chronos also provides an AMG preconditioner implementing a classical coarsening (that is, a division of the nodes into fine and coarse with no aggregation) and several interpolation schemes. Namely, classical and ExtI for Poisson-like problems and DPLS for elasticity problems. Additionally, AMG’s quality is improved by using energy minimisation [31] and its application cost reduced through prolongation filtering.
@@ 5 Experimental results
The numerical experiments consist of solving Eq. 51 using the three industrial test cases of Fig. 4: the flow around a realistic car model, a finned-tube heat exchanger, and the simulation of a wind farm. While the first two allow for exploiting reflection symmetries, the last allows for studying the further advantages of translational and rotational symmetries. All the executions rely on combined MPI and multithreaded parallelism and have been conducted on the JFF cluster at the Heat and Mass Transfer Technological Center. Its non-uniform memory access (NUMA) nodes are equipped with two Intel Xeon 6230 CPUs (20 cores, 2.1 GHz, 27.5 MB L3 cache and 140 GB/s memory bandwidth) linked to 288GB of RAM and interconnected through 7 GB/s FDR Infiniband.
@@
The numerical experiments compare AMGR- to AMG-preconditioned PCG. For simplicity, the right-hand side used in the tests is a random vector, which ensures a broad spectrum of frequencies in the resulting residual, allowing, in turn, for a complete assessment of the preconditioner’s effectiveness in damping all error components. AMG uses the standard configuration for CFD problems discussed in Section 4. That is, PMIS coarsening [18] and ExtI interpolation [17]. The smoother corresponds to a lightweight aFSAI with five nonzeros per row, and a direct method solves the coarsest level. Regarding AMGR, as discussed in Section 3.2, the first level of the hierarchy corresponds to the multigrid reduction, and subsequent levels are purely algebraic and use the same configuration as AMG. For this reason, the coarsening ratio (c-rat) and average nonzeros per row (avg nnzr) in Tables 5, 6, 7, 8, and 9 correspond to the operator at the first level of the multigrid hierarchy, i.e., after the first coarsening, which, in the case of AMGR, is induced by the reduction framework, whereas for AMG by the PMIS coarsening and ExtI interpolation. Additionally, $n_{b}$ denotes the number of subdomains in which the domain can be decomposed, i.e., the number of repeated blocks in $\bar{K}=\mathbb{I}_{n_{b}}\otimes K$. Note that $n_{b}=1$ corresponds to AMG and $n_{b}>1$ to AMGR.
As discussed in Section 3.2, the coarsening, interpolation, and smoother used in the multigrid reduction can be adapted to the problem at hand. In all the CFD problems considered, the smoother will be an aFSAI with five nonzeros per row. When it comes to the coarsening, Algorithm 2 allowed for setting the maximum interpolation distance, $k$. Table 5 illustrates the impact of $k$ on an AMGR using DPLS with energy minimisation (DPLS + EMIN). As noted in Section 3.2, the larger the $k$, the more aggressive the coarsening but the potentially more inaccurate the interpolation, leading to faster iterations but slower convergences. Notably, increasing $k$ reduces the number of coarse nodes, making the prolongation and, consequently, the reduced operator sparser. Table 5 confirms that $k=1$ retains excessive connections, as the number of iterations is practically the same as for $k=2$. On the contrary, using $k>2$ quickly degrades convergence, with $k=2$ generally being the best-performing value.
Table 5: Impact of the interpolation distance on the heat exchanger problem.
$k=1$ $k=2$ $k=3$ $k=4$
$n_{b}$ c-rat avg nnzr its c-rat avg nnzr its c-rat avg nnzr its c-rat avg nnzr its
1 0.36 14.5 20 na na na na na na na na na
2 0.36 50.8 18 0.14 37.2 19 0.07 19.9 30 0.04 12.1 907
4 0.36 50.6 17 0.15 37.7 18 0.08 21.3 27 0.05 14.7 427
8 0.37 49.5 16 0.15 37.6 18 0.09 22.6 26 0.06 16.9 283
The interpolation strategy is critical to make AMGR converge comparably to the underlying AMG. Using $k=2$ ensures a roughly twice faster coarsening but requires robust long-distance interpolations to preserve convergence. Table 6 compares the multigrid reductions arising from ExtI and DPLS with and without energy minimisation. Using $k=2$ leads to fine nodes only surrounded by fine nodes, and in such situations, ExtI is not as effective [17]. For this reason, in Table 6, it uses $k=1$. Unfortunately, neither ExtI produces an accurate interpolation nor does energy minimisation improve it, as the energy of the prolongation is already low, rendering ExtI ineffective for the multigrid reduction. On the other hand, the dynamic approach of DPLS made it robust enough to use $k=2$. DPLS was designed to accommodate multiple test vectors, which makes it perform poorly on CFD problems due to their one-dimensional near-null space [43]. However, Table 6 demonstrates how effectively energy minimisation improves the quality of DPLS interpolations while keeping the reduced operator sparse, resulting in an AMGR that converges comparably to AMG despite the aggressive coarsening. Hence, in the remainder of this section, the multigrid reduction will use $k=2$ and DPLS with energy minimisation.
Table 6: Impact of the interpolation scheme on the heat exchanger problem.
ExtI ExtI + EMIN DPLS DPLS + EMIN
$n_{b}$ avg nnzr its t-sol avg nnzr its t-sol avg nnzr its t-sol avg nnzr its t-sol
1 14.5 20 1.54 na na na na na na na na na
2 30.7 175 8.79 111.3 140 9.63 16.3 810 38.2 37.2 19 1.09
4 30.5 77 3.42 110.7 59 4.22 16.8 303 14.0 37.7 18 0.95
8 30.0 58 2.40 108.1 46 3.05 19.3 192 8.27 37.6 18 0.89
@@
The top of Fig. 4 displays the two test cases used to assess the performance of AMGR when exploiting reflection symmetries. Namely, the simulation of the flow around the DrivAer fastback car model [26] and within a finned-tube heat exchanger, on which we exploit one and three reflection symmetries, respectively. The results for the DrivAer case in Table 7 make clear the advantages of AMGR, which preserves the excellent convergence of the standard AMG and yields 43% speed-ups thanks to the combined action of the aggressive coarsening and the faster top-level smoother of Eq. 48, which leverages SpMM. The larger the number of repeated blocks in $\bar{K}=\mathbb{I}_{n_{b}}\otimes K$, the greater the benefits of replacing SpMV with SpMM. In this sense, the heat exchanger problem extends the results of Table 7 up to 3 symmetries. According to Table 8, regardless of $n_{b}$, the aggressive coarsening induced by the multigrid reduction results in almost the same coarse operator, $A_{c}$. Indeed, both the coarsening rate and density remain almost constant. In fact, when increasing $n_{b}$, the multigrid reduction becomes slightly less aggressive. Such behaviour follows from the fact that Algorithm 2 populates the interface with roughly the same inner unknowns, MIS($T^{k}$), regardless of $n_{b}$, whereas the interface grows (slightly) with $n_{b}$. However, larger $n_{b}$ implies faster top-level smoothing, resulting in up to 73% accelerations.
@@
preconditioner $n_{b}$ c-rat avg nnzr its t-sol (s) speed-up
@@
preconditioner $n_{b}$ c-rat avg nnzr its t-sol (s) speed-up
@@
AMGR 2 0.14 37.5 19 1.09 1.41
AMGR 4 0.15 37.4 18 0.95 1.62
AMGR 8 0.15 37.6 18 0.89 1.73
The results in Tables 7 and 8 make clear the potential of applying AMGR on domains arising from recurring structures, such as the $6\times 4$ wind farm in Fig. 4. Indeed, while reflection symmetries are generally restricted to $n_{b}\leq 8$, rotational and translational symmetries are not, allowing, for instance, the simulation of an arbitrarily large wind farm by just discretising a single wind turbine. Table 9 summarises the results obtained on a $6\times 4$ wind farm discretised according to [14]. The wind turbines are introduced through the immersed boundary method and, therefore, discretised with a structured grid stretched around the blades and coarsened vertically towards the atmospheric boundary layer. The first thing to note is that the optimal number of subdomains into which dividing the wind farm is not $n_{b}=24$. In fact, relatively small values of $n_{b}$ led to maximum speed-ups, as the application of AMG on the reduced operator quickly counterbalanced SpMM’s accelerations. This follows from the aforementioned balance between the effectiveness of the multigrid reduction and SpMM gains: the larger $n_{b}$, the faster SpMM, but the larger $n_{\text{ifc}}$. The wind farm problem is susceptible to such a trade-off, as the growing coarsening ratios in Table 9 indicate. Nevertheless, AMGR also proved effective on recurring structures, resulting up to 70% faster than the underlying AMG.
The quick saturation of SpMM gains as $n_{b}$ grows is due to the very high sparsity of the coefficient matrices considered, which makes the best-performing top-level smoother very lightweight, only having five nonzeros per row. Despite obtaining significant speed-ups, this compromised the advantages of AMGR, not only by making SpMM’s acceleration quickly counterbalanced by the application of AMG on the reduced operator but also by making both AMG and AMGR have comparable memory footprints. On top of that, note that the denser the matrices, the more data accesses are saved by SpMM, boosting its accelerations. In this sense, higher-order schemes or, more generally, applications entailing a denser Poisson’s equation (e.g., linear elasticity or geomechanical problems) would strengthen the benefits of SpMM and, therefore, of AMGR.
@@
preconditioner $n_{b}$ c-rat avg nnzr its t-sol (s) speed-up
@@
When it comes to the scalability of AMGR, it is worth recalling Section 4. As discussed, to replace SpMV with SpMM, it is necessary to apply a consistent domain partitioning. That is, instead of distributing the rows of $\bar{K}=\mathbb{I}_{n_{b}}\otimes K$ among all the available resources, we need to distribute the smaller $K$ so that its coefficients are effectively reused in the SpMM (analogously for the top-level smoother). Of course, distributing the smaller sub-matrix entails potentially larger communications. According to the results of Fig. 5, this is not a severe issue. On the one hand, AMGR preserves the excellent weak scalability of AMG. On the other, even if AMGR’s strong scalability is affected by the extra communication overheads, this effect is far from critical, especially considering the relatively large workloads that extreme-scale simulations typically entail.
@@ 6 Conclusions
This paper presented a multigrid reduction framework for accelerating AMG on regular domains. Given that modern supercomputers prioritise FLOP performance, most computational physics applications are memory-bound and, therefore, unable to achieve hardware’s theoretical peak throughput. To address this limitation, we first showed that given an arbitrarily complex geometry presenting reflection, translational or rotational symmetries, it is possible to apply a consistent ordering that makes the coefficient matrix (and preconditioners) satisfy a regular block structure. This, in turn, allows the standard SpMV to be replaced with the more compute-intensive SpMM, accelerating the application of the preconditioners. This paper focused on accelerating AMG owing to its numerical and computational effectiveness. However, it is worth noting that the intermediate strategies presented naturally apply to other preconditioners.
@@
AMGR, the proposed multigrid reduction framework, introduces an aggressive coarsening to the first level of the multigrid hierarchy, reducing the memory footprint, setup and application costs of the top-level smoother. While preserving the excellent convergence of AMG, replacing SpMV with SpMM yielded significant speed-ups. Remarkably enough, AMGR, the resulting preconditioner, does not have any requirement on the boundary conditions, which can be asymmetric. Hence, not only it succeeds in outperforming the standard AMG but also applies to a significantly broader range of industrial applications. Numerical experiments on industrial CFD applications demonstrated up to 70% speed-ups in the solution of Poisson’s equation with AMGR instead of AMG. Furthermore, even if AMGR potentially entails higher communication overheads, strong and weak scalability analyses revealed no serious degradation compared to AMG.
@@
The coefficient matrices arising from the incompressible CFD applications considered are remarkably sparse and, as a result, so is the best-performing top-level smoother. Indeed, best results were obtained smoothing with a very light FSAI, and even if we obtained significant speed-ups, this compromised the advantages of AMGR. Firstly, leveraging relatively few symmetries led to maximum speed-ups, as SpMM’s accelerations were quickly counterbalanced by the application of AMG on the reduced operator, whose size grows slightly with the number of subdomains. Secondly, both AMG and AMGR had a comparable memory footprint. Again, this follows from the particularly lightweight smoother required by the cases considered. In this sense, higher-order schemes or, more generally, applications entailing a denser Poisson’s equation (e.g., linear elasticity or geomechanical problems) would strengthen the advantages of SpMM and, therefore, AMGR.
@@
For all these, immediate lines of work include applying AMGR to (denser) problems arising from structural mechanics. Additionally, we plan to optimise the setup phase, which was irrelevant for the incompressible CFD applications considered, in which the coefficient matrix generally remains constant over time, making relatively large pre-processing overheads acceptable. Significant accelerations are expected from computing the top-level smoother on the base mesh instead of the entire domain. Finally, we want to tackle the GPU implementation of AMGR and study the extension of the multigrid reduction framework towards arbitrary domains, which, despite not exploiting SpMM, would benefit from the aggressive coarsening.